# Recurrent neural network wave functions for Rydberg atom arrays on kagome lattice

### v1

@@ Appendix D Numerical comparisons
$R_{b}=1.7,\delta=3.3$0.79056(1) -0.790964(5) -0.77546(1) -0.775412(4)
$R_{b}=1.95,\delta=3.3$0.59785(1) -0.59657(1) -0.56445(1) -0.56401(3)
Table 2: A table of the energies per site obtained by QMC and 2DRNN for a system size $8\times 8\times 3$ with fully periodic boundary conditions and for different values of the cutoff radius $R_{c}$. The error bars in parentheses correspond to the one-standard deviation uncertainty in the QMC and RNN energy estimates.
@@
$R_{b}=1.95,\delta=3.3$0.593828 -0.59700(4)
Table 3: A table comparing DMRG (with bond dimension $\chi=110$ for $R_{b}=1.7$ and $\chi=1000$ for $R_{b}=1.95$) and the 2DRNN ansatz in terms of the energies per site for a system size $6\times 6\times 3$ with fully periodic boundary conditions and for a cutoff radius $R_{c}=2$ and $\delta=3.3$. Note that a DMRG run on $R_{b}=1.95,\delta=3.3$ for $\chi=1500$ was not successful with a memory allocation of $120$ GB. To run our RNN simulations, an $80$ GB memory allocation was sufficient. The error bars in the parentheses correspond to the one-standard deviation uncertainty on the RNN energy.
@@ Appendix E Annealing and local minima
In Fig. 5, we demonstrate the importance of incorporating annealing in the training of RNNs applied to Rydberg atom arrays on the kagome lattice. These experiments are carried out in the highly entangled regime at $R_{b}=1.95$ and $\delta=3.3$. In panel (a), we observe that the ground state energy improves with more annealing steps $N_{a}$. We also highlight that the density saturates close to $1/6$, which raises the possibility of an odd quantum spin liquid Yan et al. (2023); Samajdar et al. (2021). However, it is not a sufficient condition as this phase is found to be the paramagnetic state according to our RNN numerics. We also outline a saturation of the second Renyi entropy to a large value in the asymptotic limit of $N_{a}$. All these numerics suggest that $N_{a}=10000$ is a good choice to converge our RNN training.
@@
Figure 5: Plots of the energy density $\langle\hat{H}\rangle/N$, density $\rho=(\sum_{i=1}^{N}n_{i})/N$ and the second Renyi entropy $S_{2}$ as function of the number of annealing steps $N_{a}$ for a system size $N=6\times 6\times 3$ and $R_{b}=1.95$ and $\delta=3.3$. We observe a saturation of all these observables in the asymptotic limit, which justifies the use of $N_{a}=10000$ in our numerical simulations in Fig. 3.

### v2

@@ Appendix D Comparison with QMC and DMRG
$R_{b}=1.7,\delta=3.3$0.790975(14) -0.790964(5) -0.77546(1) -0.775412(4)
$R_{b}=1.95,\delta=3.3$0.59785(1) -0.59666(2) -0.56445(1) -0.56401(3)
Table 2: A table of the energies per site obtained by QMC and 2DRNN for a system size 8×8×38$\times$ 8$\times$ 38 × 8 × 3 with fully periodic boundary conditions and for different values of the cutoff radius $R_{c}$. The error bars in parentheses correspond to the one-standard deviation uncertainty in the QMC and RNN energy estimates.
@@
$R_{b}=1.95,\delta=3.3$0.593828 -0.59716(2)
Table 3: A table comparing DMRG (with bond dimension $\chi=110$ for $R_{b}=1.7$ and $\chi=1000$ for $R_{b}=1.95$) and the 2DRNN ansatz in terms of the energies per site for a system size 6×6×36$\times$ 6$\times$ 36 × 6 × 3 with fully periodic boundary conditions and for a cutoff radius $R_{c}=2$ and $\delta=3.3$. Note that a DMRG run on $R_{b}=1.95,\delta=3.3$ for $\chi=1500$ was not successful with a memory allocation of 120120120 GB. To run our RNN simulations, an 808080 GB memory allocation was sufficient. The error bars in the parentheses correspond to the one-standard deviation uncertainty on the RNN energy.
@@ Appendix E Annealing and local minima
In Fig. 7, we demonstrate the importance of incorporating annealing in the training of RNNs applied to Rydberg atom arrays on the kagome lattice. These experiments are carried out in the highly entangled regime at $R_{b}=1.95$ and $\delta=3.3$. In panel (a), we observe that the ground state energy improves with more annealing steps $N_{a}$. We also highlight that the density saturates close to 1/61/61 / 6, which raises the possibility of an odd quantum spin liquid Yan et al. (2023); Samajdar et al. (2021). However, it is not a sufficient condition. This phase is found to be the paramagnetic state according to our RNN numerics. We also outline a saturation of the second Renyi entropy to a large value in the asymptotic limit of $N_{a}$. All these numerics suggest that $N_{a}=10000$ is a good choice to converge our RNN training.
Furthermore, we note that we implement annealing in QMC and report our results in Fig. 7, where we find that more annealing steps do not systematically improve QMC energies as opposed to annealing with RNNs. Finally, we highlight the strong agreement between RNN and QMC data for large $N_{a}$.
@@
Figure 7: Plots of the energy density $\langle\hat{H}\rangle/N$, density $\rho=(\sum_{i=1}^{N}n_{i})/N$ and the second Renyi entropy $S_{2}$ as function of the number of annealing steps $N_{a}$ for a system size $N=6\times 6\times 3$ and $R_{b}=1.95$, $\delta=3.3$ and $R_{c}=2$. We observe a saturation of all these observables in the asymptotic limit, which justifies the use of $N_{a}=10000$ in our numerical simulations in Fig. 3. For comparison, we also report the QMC data obtained using annealing for the energy per site and density.
Appendix F V-Score values
In Tab. 4, we report the v-score values Wu et al. (2024) for the variational calculations of our 2D RNN for $L=6,8$ and at the hard and easy points ($R_{b}=1.95$ and 1.71.71.7 respectively) for fixed $\delta=3.3$. The v-score is computed as:
[…]
where $\langle\hat{H}\rangle,\text{Var}(\hat{H})$ are respectively the expectation values and the variances of the Hamiltonian $\hat{H}$ in our optimized 2DRNN ansatz. $E_{\infty}$ is a reference energy at infinite temperature, which is computed here with a Monte Carlo estimate using $10^{5}$ samples Wu et al. (2024).
The estimated v-scores are a good indication of the quality of a variational calculation Wu et al. (2024). We find that our values of the v-score in Tab. 4 reflect the difficulty of the highly entangled paramagnetic phase at $R_{b}=1.95,\delta=3.3$ compared to the nematic phase at $R_{b}=1.7,\delta=3.3$.
Within our setup, the values of $E_{\infty}$ are about one order to magnitude bigger than $\langle\hat{H}\rangle$. In this case, the denominator of the v-score is dominated by $E_{\infty}$. To avoid this issue, we compute another metric: the energy variance per spin $\text{Var}(\hat{H})/N$. These values, reported in Tab. 4, provide a clearer distinction between the quality of $R_{b}=1.95,\delta=3.3$ variational calculation compared to $R_{b}=1.7,\delta=3.3$, where there is close to one order of magnitude difference in terms of the energy variances per spin.
$N$ Rydberg parameters V-score Energy variance per spin
6×6×36$\times$ 6$\times$ 36 × 6 × 3 $R_{b}=1.7,\delta=3.3$ $8.12\times 10^{-6}$ $1.16\times 10^{-3}$
$R_{b}=1.95,\delta=3.3$ $1.16\times 10^{-5}$ $9.14\times 10^{-3}$
8×8×38$\times$ 8$\times$ 38 × 8 × 3 $R_{b}=1.7,\delta=3.3$ $6.03\times 10^{-6}$ $8.59\times 10^{-4}$
$R_{b}=1.95,\delta=3.3$ $1.10\times 10^{-5}$ $8.70\times 10^{-3}$
Table 4: A table of V-score values and energy variances for different system sizes $N$ and different blockade radii $R_{b}$. These scores are obtained using the 2DRNN wave function trained using the hyperparameters reported in Tab. 1 for Fig. 3.
Appendix G Parameters sharing
The RNN recursion relation in Eq. (2) assumes the use of the same parameters across different sites, which is consistent with the translation invariance of the Rydberg atoms Hamiltonian. This parameter sharing property allows to iteratively train the RNN on larger system sizes by pre-training from previous system sizes.
In order to investigate the inductive bias of using weight sharing in our ansatz. We optimize our 2DRNN with no-parameter sharing, where the two-dimensional recursion relation is given as:
$\bm{h}_{i,j}$ $=f\!\Big{(}W_{ij}[\text{Neighbours}(\bm{h}_{i,j});\text{Neighbours}(\bm{\sigma}_{i,j})]+\bm{b}_{ij}\Big{)},$ (9)
using position-dependent weights and biases $W_{ij},\bm{b}_{ij}$. We also do the same for the Softmax layer, in Eq. (3), as follows:
$p_{\bm{\theta}}(\sigma_{i,j}|\sigma_{<i,j})=\text{Softmax}(U_{ij}\bm{h}_{i,j}+\bm{c}_{ij})\cdot\bm{\sigma}_{i,j},$ (10)
with site-dependent weights and biases $U_{ij},\bm{c}_{ij}$.
We optimize at 2DRNN with site-dependent parameters for a system size $N=8\times 8\times 3$ with $R_{b}=1.95,\delta=3.3$, and we report our results in Tab. 5, where we find that the RNN results are consistent with zero $q_{\rm EA}$ and zero $\gamma_{\rm RNN}$ as predicted by the 2DRNN with weight sharing. We also find that the energy per-spin $E/N$ for the 2DRNN with weight sharing is lower compared to the 2DRNN with non-weight sharing, even though the latter has $N$ more parameters. This result highlights the importance of embedding inductive bias of the Hamiltonian in the wave function ansatz to improve variational calculations.
2DRNN-NWS ($R_{c}=2$) 2DRNN ($R_{c}=2$)
$E/N$0.59447(2) -0.59666(2)
$q_{\rm EA}$ 0.001416(6) 0.000205(1)
$\gamma_{\rm RNN}$0.006(6) -0.003(20)
Table 5: A table comparing the 2DRNN results with weight sharing with the 2DRNN with no-weight sharing (2DRNN-NWS). Here we use a system size $N=8\times 8\times 3$ and $R_{b}=1.95,\delta=3.3$ for the comparison.
Appendix H Simulations runtime
To provide an idea of the simulation running time, we share a table of the times it takes to run one RNN simulation on 6×6×36$\times$ 6$\times$ 36 × 6 × 3, 8×8×38$\times$ 8$\times$ 38 × 8 × 3 lattices given the hyperparameters shared in Tab. 1. Our simulations are originally run on a P100 GPU with 12GB of memory. Our simulation data, in Tab. 6, demonstrates that we can get our RNN trained for about a day and 4 hours on a P100 GPU for $L=6$ and $L=8$, respectively. Note that we used pre-training in the $L=8$ simulation, from the $L=6$ simulation, to speed up training. We also showcase that this time can be further reduced using the A100 GPUs. Tab. 6 shows that our runtime is significantly reduced to less than 3.53.53.5 and 1.61.61.6 hours for both $L=6,8$. This runtime is expected to be further reduced with more advanced GPU resources. We also report the average runtime for our QMC simulations in Tab. 6 for different values of $R_{b}$ and at fixed $\delta=3.3$. Note that QMC run times are explicitly dependent on the Hamiltonian parameters. Our comparison confirms that RNN training times are significantly lower compared to QMC runs on a single CPU.
L RNN (P100) RNN (A100) QMC (CPU)
6 26.567 h 3.460 h 48.751 h
8 4.074 h 1.537 h 70.422 h
Table 6: Runtime (in hours) of RNNs simulation on P100 and A100 GPUs for different values of $L$ (we use $N_{\rm a}=10000$ for $L=6$ and $N_{\rm conv}=10000$ converge steps for $L=8$). We also report the average QMC runtime on a single CPU for different values of $R_{b}=1.95$ and at fixed $\delta=3.3$. All of these simulations are conducted with a cutoff radius $R_{c}=2$. The lowest values are highlighted in bold.